the vapour density of a monoatomic gas is 20. what is the volume occupied by 20grams of this gas at ntp?

Grade:11

1 Answers

Mayank Dixit
6 years ago
molecular mass of gas = 2*V.D. =2*20 =40no of moles of gas = weight/ molecular mass = 20/40vol oF gas = no of moles * V = 20/40* 20 = 10